KClO4 can be prepared by Cl2 and KOH by a series of reactions as given below :Cl2 + KOH ----> KCl + KClO + H2O ​KClO ----> KCl + KClO3KClO3 ----> KClO4 + KCl​ calculate mass of Cl2 in gm required to produce 1385 gm KClO4calculate the total mass of KCl produced in 1st, 2nd and 3rd reaction

Grade:11

2 Answers

Vikas TU
6 years ago
The arrangement of conditions includes; 
Cl2 + 2KOH → KCl + KClO + H2O - - (1) 
3KClO → 2KCl + KClO3 - - (2) 
4KClO3 → 3KClO4 + KCl - - (3) 
From the eqn(3) 
The proportion KClO4 :KCl = 3:1 
Implying that we have 3 moles of KClO3 We have 1mole of KCl 
Since the molar mass of KClO4 is 138.55, we have 156/138.55 
Moles delivered = 1.126 moles 
On the off chance that we accept 100 % yield these originated from (1.125 *4/3) moles of KClO3 
=1.501 moles of KClO3 required to deliver 156g of KClO4 
Moles of KClO expected to deliver 1.501 moles of KClO3 
= 1.501 * 3 =4.504 moles of KClO from eqn(2) 
From eqn(1) 
Moles of Cl2 = moles of KClo =4.504moles. 
Along these lines, the mass of Cl2 = 4.504 * 71 
Mass of Cl2=319g to deliver 156g of KClO4 if 100% yield is accepted in the three compound responses.
Baba
2 years ago
i) Let mole of Cl2
 required = x
Cl2+ 2KOH−> KCl + KClO + H2O
            (X)                      (x)      (x) 
 
3KClO −> 2KCl + KClO
x               2x/3      x/3
4KClO3−>3KClO4 + KCl
 x/3            x/4               x/12
Mole of KClO4 formed = 1385/138.5(molar weight) 
=10
 
x/4 =10 , x = 40
weight of Cl2 required for the reaction = 40 × 71 = 2840 gm
 
(ii) KCl produced from Ist, IInd and IIIrd reaction = x+(2x/3) +(x/12) moles
=21x/12 mole=(21/12) ×40=70 mole
weight of KCl produced =70 × 74.5 = 5215 gram